OpenStudy (anonymous):

Which of the following is the equation of g(x) if the parent function is f(x) = x2 and a = 1/4, h = -2, and k = -10?

OpenStudy (anonymous):

(a) g(x) = -1/4(x + 2)^2 - 10 (b) g(x) = 1/4(x - 2)^2 + 10 (c) g(x) = 1/4(x - 2)^2 - 10 (d) g(x) = 1/4x(x - 2)^2 - 10

OpenStudy (anonymous):

are you gonna belive me and not act like im just putting out any old answer?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes I believe you, but I need to show my work so if I don't understand how to get the answer...

OpenStudy (anonymous):

I will have an exam on this later, I don't want to be beholden to other ppl in order to get the right answer

OpenStudy (anonymous):

@Abmon98 @nincompoop

OpenStudy (anonymous):

go on mathway.com and type your question it will give you the work AND the answer:)